For Monday mornings staff meeting, Jim bought 2 bags of bagels and 3 packages of cream cheese and paid $9.75 (excluding sales tax). For Fridays meeting, he bought 4 bags of bagels and 2 packages of cream cheese and paid $12.50 (again, excluding sales tax). How much do bags of bagels and packages of cream cheese cost? : For Monday mornings staff meeting, Jim bought 2 bags of bagels and 3 packages of cream cheese and paid $9.75 (excluding sales tax). For Fridays meeting, he bought 4 bags of bagels and 2 packages of cream cheese and paid $12.50 (again, excluding sales tax). How much do bags of bagels and packages of cream cheese cost?
Answer by mangopeeler07(441)

x=cost of bagels (in cents)
y=cost of cream cheese (in cents)
coefficient=quantity

Jim bought 2 bags of bagels and 3 packages of cream cheese and paid $9.75:
2x+3y=975

He bought 4 bags of bagels and 2 packages of cream cheese and paid $12.50:
4x+2y=1250

So you have two equations:
2x+3y=975
4x+2y=1250

Multiply the first equation by negative two:
-4x-6y=-1950
4x+2y=1250

Add the equations
-4y=-700

You have successfully eliminated x from the equation (for now).

Divide by -4
y=175

That is the cost of the cream cheese in cents. Plug that back into either original equation, how about the first one:
2x+3(175)=975

Simplify
2x+525=975

Subtract 525 from both sides
2x=450

Divide by 2
x=225

That is the cost of the bagels in cents.
So,

1 Package of Cream cheese=$1.75
1 Bag of Bagels=$2.25